If Rebecca were three times as old as she was five years ago, she will be twenty-one more than two times her current age. How old is Rebecca?
Answer by ankor@dixie-net.com(22740)

If Rebecca were three times as old as she was five years ago, she will be twenty-one more than two times her current age. How old is Rebecca?
:
Let x = her current age
then
(x-5) = her age 5 yrs ago
:
The equation:
3(x-5) = 2x + 21
3x - 15 = 2x + 21
3x - 2x = 21 + 15
x = 36 yrs old now
:
:
Check solution in the given statement:
3(36-5) = 2(36) + 21
3 * 31 = 72 + 21